OpenStudy (anonymous):

.simplify the expression -84/14 -6 6 98 -70

OpenStudy (solomonzelman):

\[\huge\color{blue}{ \frac{-84}{14} } \]

OpenStudy (solomonzelman):

\[-84=14 \times 6\]

OpenStudy (anonymous):

whats the god damn rule that's all I need is it positive or negative

OpenStudy (solomonzelman):

\[\color{blue}{ \frac{-84}{14}~~~->~~~\color{blue}{ \frac{(-6) \times 14}{14}}~~~->~~~\color{blue}{ \frac{(-6) \times \cancel{14 } }{\cancel{14 }}~~~=?}}\]

OpenStudy (solomonzelman):

can you tell me the answer now, after I canceled the 14 on top and bottom for you?

OpenStudy (anonymous):

why r you fluttering with me why don't you flutter with someone else im not a pissing post

OpenStudy (solomonzelman):

\[\color{blue}{ \frac{28}{4} ~~->~~~\frac{7 \times 4}{4} ~~->~~~\frac{7 \times \cancel{ 4}}{\cancel{ 4}} =7}\]

OpenStudy (solomonzelman):

@lakotareid1232 can you do your example looking at my replies?

OpenStudy (solomonzelman):

I meant your question not your example sorry.

OpenStudy (anonymous):

no I no its 6 but is it positive or negative

OpenStudy (solomonzelman):

what do you think?

OpenStudy (anonymous):

divide -84 and 14 and thats ur answer

OpenStudy (the_fizicx99):

That's rude, a negative and a positive will = a negative..

OpenStudy (solomonzelman):

What's rude? was I being rude?! HOW?

OpenStudy (the_fizicx99):

Not you, him.

OpenStudy (solomonzelman):

Solmyr?

OpenStudy (anonymous):

thank you @tHe_FiZiCx99

OpenStudy (the_fizicx99):

._. oh god U_U the asker..

OpenStudy (solomonzelman):

Oh, okay.

OpenStudy (anonymous):

how am I being rude @tHe_FiZiCx99

OpenStudy (the_fizicx99):

Well you wrote "why r you fluttering with me why don't you flutter with someone else im not a pissing post " >.<

OpenStudy (anonymous):

-(84/14) = -6

OpenStudy (solomonzelman):

@lakotareid1232 \(\color{red}{I~~~know~~how~~much~~knowledge~~you ~~really~~have,~~if~~even~~after}\)\(\color{red}{my~~help~~you~~can't~~do~~it.~~I~~made~~it~~as~~easy~~as~~possible,~~you}\) \(\color{red}{have~~to~~just ~~swallow,~~I~~chewed~~it~~all~~for~~you!}\) You were RUDE at the begging, and you know how, but that's not the main problem, the problem, is that you don't want to learn, you are LAZY! I don't care if you are rude, I would like a rude but not lazy as opposed to not rude and lazy, and certainly not rude and lazy.

OpenStudy (solomonzelman):

What will you do i Trigonometry, calculus... not every problem you ask will be answered like this, you know?

OpenStudy (the_fizicx99):

Lol Solo, :D you go boo

OpenStudy (anonymous):

im not lazy I knew the answer I just didn't know if it was positive or negative

OpenStudy (solomonzelman):

ty, I just want this kid to learn something

OpenStudy (solomonzelman):

@lakotareid1232 "I knew the answer I just didn't know if it was positive or negative " you realize that makes NO SENSE what so ever, because if you don't know if it is positive or negative then YOU DON"T KNOW the answer.

OpenStudy (anonymous):

I knew it was 6

OpenStudy (the_fizicx99):

It's not 6.. >.>

OpenStudy (anonymous):

how did u come to the conclusion of thinking it was 6??

OpenStudy (anonymous):

because u can just come up out of the blue and say i think its 6

OpenStudy (anonymous):

cant*

OpenStudy (solomonzelman):

He didn't come to any conclusions, sorry I have to say this but you are LAZY-GUESSER, I don't see any point wasting my tie and time of other people on you, because you don't even want to learn. GUYS IGNORE REPLIES FROM THIS POST!

OpenStudy (the_fizicx99):

84/14 is 6, however they asked for -84/14, A negative and a positive is a negative. So based of that, which is it

OpenStudy (anonymous):

-6 all I needed was the rule not the answer im homeschooled so I cant just ask my teacher for help

OpenStudy (the_fizicx99):

Well there's your rule :3

OpenStudy (anonymous):

what?